OpenStudy (anonymous):

The base of a triangle exceeds its height by 17 inches. If its area is 55 square inches, find the base and height of the triangle.

OpenStudy (anonymous):

Hello all.

OpenStudy (anonymous):

sup

OpenStudy (anonymous):

Any ideas?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

I thought I should set up the problem this way: 55 = 1/2(h+h+17)

OpenStudy (anonymous):

But then I got the wrong answer, as usual

OpenStudy (mathstudent55):

let h = height, then let b = base = h + 17 Area of triangle: A = (bh)/2 55 = (h + 17)h/2

OpenStudy (anonymous):

Right, exactly..so 55 = 2h+17/2, and thus h+17/2 = 55. Then, 55*2 = h+17

OpenStudy (mathstudent55):

Your're close, but the formula for the area of a triangle is the base times the height over 2, not the base plus the height over 2.

OpenStudy (helder_edwin):

right \[ \large 55=\frac{(h+17)\cdot h}{2} \]

OpenStudy (anonymous):

Um...but I already multiplied the b by the h. It's just that here, b = h + 17

OpenStudy (mathstudent55):

it's not (2h + 17)/2 = 55, it's (h^2 + 17h)/2 = 55

OpenStudy (helder_edwin):

use the distributive law on the denominator

OpenStudy (helder_edwin):

\[ \large 55=\frac{h^2+17h}{2} \]

OpenStudy (anonymous):

Okay. So you end up with h^2 + 17h = 110

OpenStudy (anonymous):

Then...h^2+17h-110

OpenStudy (mathstudent55):

Right. Now subtract the 110 to the left and solve the quadratic eq.

OpenStudy (mathstudent55):

Set it equal to zero and solve.

OpenStudy (anonymous):

(x+22)(x-5) = 0, x = -22, 5

OpenStudy (anonymous):

Sorry...I meant h. So h = 5, and then b = (5) + 17 = 22

OpenStudy (anonymous):

Thank you for all your help!!

OpenStudy (mathstudent55):

Not x, but h. Now you have h = -22 or h = 5. A negative number is meaningless as the height of a triangle, so dicard it. yes, you got it.

OpenStudy (mathstudent55):

wlcm